# A Simple Baseline for Stable and Plastic Neural Networks

### v1

@@ 1 Introduction
Some recent efforts attempt to address this by leveraging multi-scale feature encodings from pre-trained backbones, structure-wise distillation losses to preserve hierarchical representations, and bespoke stability–plasticity normalization layers to balance adaptation and retention  Jung et al. (2023); Elsayed & Mahmood (2024). While effective, these solutions incur significant architectural complexity, rely on auxiliary distillation objectives, and depart from the simplicity desired for a baseline framework. In this work, we explore a simple yet effective approach to continual learning that preserves high plasticity without undermining stable representations or imposing heavy computational costs. Specifically, we introduce:
@@ 2 Related Work
Continual learning, also referred to as lifelong learning Parisi et al. (2019), requires models to learn from a sequence of non-stationary data streams while managing both network plasticity and stability Kim & Han (2023). Traditional neural networks, when trained sequentially on multiple tasks, often succumb to catastrophic forgetting Riemer et al. (2018), erasing earlier representations, or experience diminished adaptability to new tasks.
@@
Maintaining plasticity ensures that models can adapt their parameters in response to shifts in data distributions Berariu et al. (2021) Two main phenomena impact the network’s plasticity: first, shifts in the pre-activation distribution during training can cause neurons to become “dead” Lin et al. (2015), “dormant” Sokar et al. (2023), or effectively linearized Lyle et al. (2024), reducing the network’s expressive capacity; second, unchecked growth in parameter norms sharpens the loss landscape, driving activation and normalization layers toward saturation and diminishing sensitivity to gradient updates Damian et al. (2022). To counteract these effects, three main categories of methods have emerged. One line of work periodically resets or overwrites selected weights or layers to increase the network’s ability to learn new patterns Dohare et al. (2024); He & Liu (2025). A second approach focuses on activation-level interventions using adaptive activation functions or carefully calibrated noise injections to preserve plasticity without altering the underlying training dynamics Delfosse et al. (2021); Kuo et al. (2021). Finally, architectural expansion methods introduce additional modules or subnetworks that specialize in handling new tasks, which allows the base model to retain existing representations intact Liu et al. (2025).
@@
Maintaining stability is essential not only to prevent catastrophic forgetting but also to facilitate efficient transfer to related tasks Weiss et al. (2016); Zhuang et al. (2020). One common strategy is to retain and replay a small set of representative examples from past tasks, thereby reinforcing old knowledge whenever the network trains on new data Shin et al. (2017). Other methods dynamically adjust the network’s structure, either by dedicating subnetworks to each task or by growing new modules on demand, to protect previously learned features Yoon et al. (2017). Regularization-based approaches, such as Elastic Weight Consolidation, impose penalties on updates to parameters deemed important for earlier tasks, using measures like the Fisher information matrix to identify and protect critical weights Schwarz et al. (2018).
@@
Recent works tackle the stability-plasticity dilemma in various contexts through task sequencing, alternating update schemes, and transfer learning. For example, Jaziri et al. (2024) propose subspace expansion which alternates between weight regularization for better stability and adaptive activations for higher plasticity in a curriculum learning setting. Scialom et al. (2022) continual fine-tuning of pre-trained models refines large-scale features without full retraining. Integrated frameworks further combine regularization, distillation, and specialized normalization to balance adaptation and retention Jung et al. (2023). Yet, these methods often add architectural or computational overhead, underscoring the need for a simple baseline like RDBP proposed in this work.
@@ 3.1 Maintaining Plasticity: ReLUDown
This issue can be effectively addressed by introducing a linear component with a non-zero gradient for the range from the hinge point $d<0$ to $\infty$- ∞. This approach preserves the non-linearity of the activation function while ensuring non-zero gradients on both sides of the zero-encoding, thereby enabling the model to implicitly stabilize its preactivation distribution by the standard training process. The activation function can be formulated as follows:
@@ 3.2 Maintaining Stability: DecreaseBackpropagation
$\text{bp}_{decrease}(n,l,f,a)=\text{bp}_{standard}*(1-(l*f)+(l*f)*a^{-n})$
@@ 4.2 Results and Discussion
In Figure 2, we illustrate the impact of various activation functions on the neural network’s ability to maintain plasticity (left panel). The results demonstrate that the network utilizing the ReLU activation function exhibits a decline in performance, despite similar task complexity. In contrast, networks employing the ReLUDown and PAU activation functions are able to maintain their plasticity over time. Additionally, our proposed activation function achieves this while also reducing training time by approximately 20% compared to Padé Activation Units (see Appendix A.5.3).
@@ A.2 Example Images
Figure 5: Example Images of the downsampled ImageNet Dataset. : Baby, Dog, Bridge, Bird
@@ A.4 Decrease Backpropagation
Figure 7: Gradient Factor for the different Layers of the Decrease Backpropagation algorithm with a Speed Factor of $a=1.005$ and a Decrease Factor of $f=0.15$ based on the task.

### v2

@@ 1 Introduction
Some recent efforts have attempted to address this by leveraging multi-scale feature encodings from pre-trained backbones, structure-wise distillation losses to preserve hierarchical representations, and bespoke stability–plasticity normalization layers to balance adaptation and retention  Jung et al. (2023); Elsayed & Mahmood (2024). These methods have demonstrated notable performance improvements across various continual learning benchmarks, showcasing their ability to mitigate forgetting while enabling efficient task adaptation. However, despite their effectiveness, these solutions incur significant architectural complexity, rely on auxiliary distillation objectives, and depart from the simplicity desired for a baseline framework. In this work, we explore a simple yet effective approach to continual learning that preserves high plasticity without undermining stable representations or imposing heavy computational costs. Specifically, we introduce:
@@ 2 Related Work
Continual learning, also referred to as lifelong learning  Parisi et al. (2019), requires models to learn from a sequence of non-stationary data streams while jointly addressing both network plasticity and stability  Kim & Han (2023). Traditional neural networks, when trained sequentially on multiple tasks, often experience catastrophic forgetting  Riemer et al. (2018), where previously learned representations are overwritten, and their ability to adapt effectively to new tasks is significantly reduced.
@@
Maintaining plasticity ensures that models can adapt their parameters in response to a change in the data distribution Berariu et al. (2021). Two main phenomena impact the network’s plasticity: first, shifts in the pre-activation distribution during training can cause neurons to become “dead” Lin et al. (2015), “dormant” Sokar et al. (2023), or effectively linearized Lyle et al. (2024), reducing the network’s expressive capacity; second, unchecked growth in parameter norms sharpens the loss landscape, driving activation and normalization layers toward saturation and diminishing sensitivity to gradient updates Damian et al. (2022). To counteract these effects, three main categories of methods have emerged. One line of work periodically resets or overwrites selected weights or layers to increase the network’s ability to learn new patterns Dohare et al. (2024); He & Liu (2025). A second approach focuses on activation-level interventions using adaptive activation functions or carefully calibrated noise injections to preserve plasticity without altering the underlying training dynamics Delfosse et al. (2021); Kuo et al. (2021). Finally, architectural expansion methods introduce additional modules or subnetworks that specialize in handling new tasks, which allows the base model to retain existing representations intact Liu et al. (2025).
@@
Maintaining stability is essential not only to prevent catastrophic forgetting but also to facilitate efficient transfer to related tasks  Weiss et al. (2016); Zhuang et al. (2020). One common strategy is to retain and replay a small set of representative examples from past tasks, thereby reinforcing old knowledge whenever the network trains on newly introduced data  Shin et al. (2017). Other methods dynamically adjust the structure of the network, either by dedicating subnetworks to each task or by growing new modules on demand, to protect previously learned features  Yoon et al. (2017). Regularization-based approaches, such as Elastic Weight Consolidation, impose penalties on updates to parameters deemed important for earlier tasks, using measures like the Fisher information matrix to identify and protect critical weights Schwarz et al. (2018).
@@
Recent works tackle the stability-plasticity dilemma in various contexts through task sequencing, alternating update schemes, and transfer learning. For example, Jaziri et al. (2024) propose multiple subspace expansions during training, which alternate between weight regularization for better stability and adaptive activations for higher plasticity in a curriculum learning setting. While Scialom et al. (2022) employs continual fine-tuning of pre-trained models, refining large-scale features without full retraining. Integrated frameworks further combine regularization, distillation, and specialized normalization to balance adaptation and retention Jung et al. (2023). Yet, these methods often add architectural or computational overhead, underscoring the need for a simple baseline like RDBP proposed in this work.
@@ 3.1 Maintaining Plasticity: ReLUDown
This issue can be effectively addressed by introducing a linear component with a non-zero gradient for the range from the hinge point $d<0$ to $\infty$- ∞. This approach preserves the non-linearity of the activation function while ensuring non-zero gradients on both sides of the zero-encoding, thereby enabling the model to implicitly stabilize its preactivation distribution by the standard training process. The activation function is displayed in the appendix and can be formulated as follows:
@@ 3.2 Maintaining Stability: DecreaseBackpropagation
$\text{bp}_{decrease}(n,l,f,a)=\text{bp}_{standard}\times(1-(l\times f)+(l\times f)\times a^{-n})$
@@ 4.2 Results and Discussion
In Figure 2, we illustrate the impact of various activation functions on the neural network’s ability to maintain plasticity (left panel). The results demonstrate that the network utilizing the ReLU activation function exhibits a decline in performance, despite similar task complexity. In contrast, networks employing the ReLUDown or Padé Activation UnitDelfosse et al. (2021) activation functions are able to maintain their plasticity over time. Additionally, our proposed activation function achieves this while also reducing training time by approximately 20% compared to the Padé Activation Units (PAU)(see Appendix A.5.3).
@@ 5 Conclusion and Future Work
Acknowledgments
This work was supported by the KIBA Project (Künstliche Intelligenz und diskrete Beladeoptimierungsmodelle zur Auslastungssteigerung im Kombinierten Verkehr - KIBA) under reference number 45KI16E051, funded by the German Ministry of Transportation.
@@ A.2 Example Images
Figure 5: Example Images of the downsampled ImageNet Dataset. : Baby, Dog, Bridge, Bird, Oven
@@ A.4 Decrease Backpropagation
Figure 7: Gradient Factor for the different Layers of the DBP algorithm with a Speed Factor of $a=1.005$ and a Decrease Factor of $f=0.15$ based on the task.